OpenStudy (anonymous):

FAN AND MEDAL**PLEASE HELP ME ITS MY LAST QUESTION, (i dont understand the fk qs) :(

OpenStudy (anonymous):

Now that the curve pieces are determined, use those pieces as sections of a complete coaster. By hand or by using a drawing program, sketch a design of Ray and Kelsey's coaster that includes the shape of the g(x) and f(x) functions that you chose in the Parts A and B. You do not have to include the coordinate plane. You may arrange the functions in any order you choose, but label each section of the graph with the corresponding function for your instructor to view.

OpenStudy (anonymous):

@igreen @Peaches15 @ganeshie8

OpenStudy (anonymous):

@EmilyF03 This will help: http://www.peeranswer.com/study/mathematics/55b2d331b0447d5d5d4ddbea

OpenStudy (anonymous):

Here @EmilyF03 https://answers.yahoo.com/question/index?qid=20150725104721AAViLwr

OpenStudy (anonymous):

@Michele_Laino @johnweldon1993

OpenStudy (anonymous):

uuugghhh you stole my answer i was going to do that

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

@pooja195

OpenStudy (michele_laino):

I need to know the equations of the function f(x) and g(x)

OpenStudy (anonymous):

okay wait

OpenStudy (anonymous):

f(x) = ax2 + bx + c. g(x) = x3 − x2 − 4x + 4

OpenStudy (anonymous):

@Michele_Laino help plz

OpenStudy (michele_laino):

do you know the values of the coefficients a, b, and c, of f(x) function?

OpenStudy (anonymous):

no ;/

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

(a, 0), (b, 0), and (c, 0)?

OpenStudy (anonymous):

that doesnt help

OpenStudy (anonymous):

i can sort of visualize it as something like this